OpenStudy (anonymous):

simplify (x^5/8) / (x^1/6)

OpenStudy (anonymous):

subtract the exponents

OpenStudy (anonymous):

\[\frac{5}{8}-\frac{1}{6}=\frac{11}{24}\]

OpenStudy (anonymous):

so answer is \[x^{\frac{11}{24}}\]